OpenStudy (anonymous):

2. Graph the function Draw a chart to show at least three ordered pairs that are on the graph

OpenStudy (anonymous):

The eques\[Y=x ^{2} +2x-3\]

OpenStudy (anonymous):

That's the equation for it.

OpenStudy (anonymous):

But I see that this cannot be answered because there is not online whiteboard

OpenStudy (anonymous):

OK, here is an example y = 2x 0 = 2*0 1 = 2*0.5 2 = 2*1 etc etc so pairs are 0,0 0.5,1 1,2

OpenStudy (anonymous):

The bottom is the points that would be plotted?

OpenStudy (anonymous):

Yes, they are obtained from the top....

OpenStudy (precal):